### v2

@@ 1 Data Bias: Train-Serving Skew
The goal of classical ML research and statistical learning theory is to achieve stable performance in an i.i.d. environment, the training data and the unseen data come from the same distribution bishop2006pattern . However, the iid assumption is not satisfied in most cases because the construction of training datasets with high probability to represent the true data distribution (the inputs in production) is extremely difficult sugiyama2012machine . Real-world data is multi-faceted and virtually infinite, whereas training datasets are finite and constrained by the resources available during the dataset preparation. There are many ways that data bias can occur and misrepresent training datasets for real-world applications. This leads to a common failure mode known as the train-serving skew: models that perform well in development but poorly in deployment.
@@
There are two main categories of data bias in ML applications: erroneous bias and discriminatory. Erroneous bias can be viewed as a systematic error caused by faulty assumptions. For instance, due to selection bias marlin2012collaborative or sampling bias mehrabi2021survey , the chosen training samples may not be able to represent the real data distribution. Measurement bias happens when the device used to measure the signal has systematic value distortion that tends to skew the data in a particular direction that prevents the generalization of other data collected by other devices. Discriminatory bias is one of the concerns in AI nondiscrimination and fairness. As opposed to fairness, discriminatory bias reflects an algorithm’s unfair behaviors toward a certain group or individual, such as producing discriminatory content or performing less well for some people shah2019predictive . The model is likely to inherit this discrimination bias and leads to undesirable performance on the minority groups in real-world settings.
@@
A robust ML workflow requires dealing with data bias, especially addressing the long-tailed distribution modeling and the edge cases. Indeed, the long-tailed distributions of data are extremely common in machine learning, reflecting the state of the real world and typical data collection practices. These distributions are influenced by natural events, and not necessarily introduced by skewed data collection. Nonetheless, ML algorithms fail to handle them properly as they are statistically optimized to perform well on common inputs (i.e. the head of the distribution) but struggle where examples are sparse (the tail). The tail often comprises the largest proportion of possible inputs, which makes their inclusion a laborious iterative procedure, i.e., collecting new data and retraining to account for edge cases. In safety-critical applications, an ML model that performs well on most cases but fails on infrequent edge cases might not be usable because these failures could cause catastrophic consequences. For this reason, major self-driving car companies strive to gather edge cases brooks2017edge , this is also applicable to safety-critical applications such as medical diagnosis, and traffic control.
@@ 2 Model Complexity: A Double-Edge Sword
Conventionally, it has been assumed that the use of models with increasing capacity will systematically result in overfitting the training data. Hence, the capacity of the models is usually controlled either by limiting the size of the model (number of parameters) or via various explicit or implicit regularizations, such as early stopping yao2007early , batch normalization ioffe2015batch , dropout srivastava2014dropout , and weight decay krogh1991simple . This aims to push learning to a subspace of a hypothesis with manageable complexity and reduce overfitting zhang2021understanding . Nonetheless, researchers have found that increasing model complexity not only allows for perfect interpolation but also results in low generalization error. Various studies have been conducted to analyze such overparameterized models, i.e., trainable parameters are much higher in number than the training data points. From the statistical viewpoint, the majority of overparameterized models exhibit a double-descent effect belkin2018understand ; mei2022generalization . In fact, the generalization error follows the traditional U-shaped curve until a specific point, after which the error decreases, and reaches a global minimum in the overparameterized regime. According to the double-descent phenomenon, the minimum generalization error tends to appear at infinite complexity, i.e., the more overparameterized the model, the smaller the error. From the optimization viewpoint, the good generalization behavior of highly overparameterized models is also commonly attributed to the inductive bias of gradient-based algorithms which helps with selecting models that generalize well despite the non-convexity, e.g., bartlett2021deep . Intuitively, the large number of hidden units here represent all possible features, and hence the optimization problem involves just picking the right features that will minimize the training loss. This suggests that as we over-parametrize the networks, the optimization algorithms need to do less work in tuning the weights of the hidden units to find the right solution.
@@
A disadvantage of overparameterized deep learning architectures is that they are highly susceptible to perturbations in adversarial or non-adversarial settings, compared to conventional, less sophisticated models. To illustrate the inherent brittleness of overparameterized neural networks, we refer to the notion of “neuron coverage” that is inspired by the code coverage in traditional software testing deepxplore . It involves generating synthetic test input data to trigger the neurons that have not been activated by the original test data. The success of this coverage criterion suggests that only a subset of the parameters is responsible for capturing the patterns needed for the task. The rest of the parameters might be unoptimized (almost stalled at initial random weights) or have received fewer updates over the training. The presence of these suboptimal subnetworks might not harm the model’s performance under iid conditions. However, it does affect the model’s robustness negatively. Any unusual changes in inputs that accidentally activate these neurons could lead to unpredictable model behaviors. In addition, any attacker can exploit the larger space of these suboptimal neurons by designing a malicious input that yields a particular model’s output.
@@ 3 Underspecified ML Pipeline: One Pipeline, Many Models
In order to solve an ML problem, we expect the model to encode some essential structure of the underlying distribution, which is inferred from the data using a designed ML pipeline, and is often what makes a model credible. Nevertheless, many explorations of the failures of ML pipelines that optimize for iid generalization, reveal a conflict between iid generalization and encoding credible inductive biases. It is called structural failure mode, as it is often diagnosed as a misalignment between the predictor learned by empirical risk minimization and the causal structure of the desired predictor arjovsky2019invariant . In medical applications of ML, training inputs often include markers representing a doctor’s diagnostic judgments oakden2020hidden . An analysis of a CNN model for diagnosing skin lesions showed that it relied heavily on surgical ink markings around skin lesions that doctors had deemed cancerous winkler2019association . In these situations, a predictor with credible inductive biases cannot achieve optimal iid generalization in the training distribution, because there are so-called “spurious” features that are strongly associated with the label in the training data, but are not associated with the label in practically important settings. In fact, an iid-optimal predictor would incorporate the ink markings as they are highly correlated with positive cases, but these markings would not be expected to be present in deployment, where the predictor would itself be part of the workflow for making a diagnostic judgment. There is clearly an underspecification problem as there is not enough information (a lack of positivity) in the training distribution to distinguish between credible inductive biases and spurious relationships. Geirhos et al. shortcutLearning connect this underspecification issue to the notion of “shortcut learning”. They point out that there may be many predictors that generalize well in iid settings, but only a few of them align with the intended solution to the learning problem. Shortcut learning resembles surface learning of students in classrooms, relying on simple decision rules to pass an exam shortcutLearning . The problem with shortcuts is that they might go unnoticed during the iid testing and only occur in deployment scenarios. In the absence of large, diverse datasets, their risk is higher. For example, diagnostic data for rare or novel diseases is usually limited to small datasets, and unbiased validation data could be difficult to acquire.
@@
Modern ML pipelines are poorly set up for satisfying the system requirements underspecificationMLmodel . Their iid evaluation procedure often results in multiple models with equivalent (similar) predictive risk (performance) while they encode substantially different inductive biases. This implies that the ML pipeline could not distinguish between these iid-optimal models despite their potential differences in terms of robustness. The ML pipelines must be specified and designed in a way that promotes the selection of the model, encoding credible inductive biases, to bridge the gap between testing behavior and deployment behavior.
@@ 1 Adversarial Attacks: Categories and Aims
The first proposed evaluation methods for ML model robustness are the adversarial attacks, where the community invents many ways to carefully craft perturbations that can deceive a given ML model. The purpose of such attacks is to produce an adversarial example (AX), i.e., an input $x^{\prime}$ close to a valid input $x$ according to some distance metric (i.e., similarity) or admissibility criteria (i.e., semantically-preserving modification ranges), whose model’s predictions, denoted as $f(x^{\prime})$ and $f(x)$, respectively, are different ($f(x^{\prime})\neq f(x)$). First, they can be categorized into white-box attacks and black-box attacks according to how much knowledge an attacker has about the subject model. White-box attacks are implemented with direct access to the model or its training data, whereas black-box attacks can only access the target model through queries: pairs of inputs and outputs. The growth of black-box attacks started with the discovery of the transferability of adversarial examples papernot2016transferability . Cross-model transferable, where attackers can construct adversarial examples in known deep learning models and subsequently attack a related unknown model. Cross training-set transferable refers to the attacks that exploit shared vulnerabilities across different datasets or domains. Second, adversarial attacks can be targeted or non-targeted. If we consider the example of image classification, targeted attacks aim to force the classifier to output a particular (chosen) class, whereas untargeted attacks attempt to make it return any class other than the original label. This categorization of adversarial attacks is based on the attacker’s goal and information access, which determines the relevance of an attack for an application. For instance, targeted white-box attacks are more suitable for security concerns, while untargeted black-box attacks can be appropriate for assessing robustness to noise.
@@
The most notorious adversarial attack is the Fast Gradient Sign Method (FGSM) FGSM . It operates as a one-step method by computing the gradient of the model’s cost function ($J$) with respect to the input data ($x$) and then perturbing the input along this gradient direction. Specifically, it alters the input ($x$) by adding noise ($\eta$) in the direction that maximizes the loss, using a magnitude defined by an epsilon value ($\epsilon$) to limit the perturbation within a certain range. The attack aims to maximize the loss by perturbing the input data without exceeding the epsilon-boundary, calculated using a distance metric (e.g., $L_{\infty}$ or $L_{2}$ norm). The formulation can be represented as:
@@
Iterative-FGSM or IFGSM IFGSM is a straightforward enhancement to FGSM. It involves iteratively applying the same step as FGSM with a small step size and clipping the pixel values of intermediate results after each step to ensure proximity to an $\epsilon$-neighbourhood of the original input. The attacking scheme can be represented mathematically as follows:
[…]
Projected Gradient Descent (PGD) PGD attack represents one of the most established and efficient methods.It is a variant of Iterative FGSM, differing in their initialization strategy. More specifically, PGD initializes the example randomly within the sphere of interest determined by the $L_{\infty}$ norm and does random restarts, whereas IFGSM starts from the original point.
[…]
Papernot et al. JSM introduced a jacobian-based class of attacks to craft adversarial examples based on computing forward derivatives. Knowing the network architecture, an adversary can construct saliency maps that identify features of the input that most significantly impact output classification. Indeed, saliency maps simonyan2013deep were originally designed for visually explaining the input features the classifier used to generate its output decisions. The map indicates how much influence each input feature (e.g. pixel) has on predicting a specific class. By increasing a few high-saliency pixels, $x(i)$, an attacker can influence the model to predict the adversarial class, $t$, that does not match the true class label, $y$, of a given input $x$, and this may result in a misclassification of the input.
DeepFool deepfool algorithm iterates to find the smallest perturbation that could misclassify an input. It represents the classifier’s decision space with linear hyperplane boundaries and iteratively shifts the image towards the closest boundary until misclassification. The algorithm takes an input $x$ and a classifier $f$, outputs the minimal perturbation, initializes the adversarial input with the original, and loops while the true label matches the adversarial label. It calculates the minimal perturbation by projecting the input onto the closest hyperplane, adds it to the input, and tests for misclassification. Finally, it returns the minimal perturbation.
Xiao et al. AdvGAN propose to generate adversarial examples with generative adversarial networks (GANs), which can learn and approximate the distribution of original instances. Using GAN-style learning, they train the generator to create perturbations that deceive the target model, while the discriminator is optimized to differentiate between original and perturbed instances. Once the generator is trained, it can generate adversarial perturbations efficiently for any instance to derive a visually-realistic crafted version that will likely fool the target model.
@@
Black-box adversarial attacks are less effective than white-box approaches because they often require a higher number of queries, but they can expose different adversarial examples and are better representative of external attack system simulations. Narodytska et al. narodytska2017simple performed a local-search-based attack. Chen et al. chen2017zoo and Bhagoji et al. bhagoji2018practical proposed black-box attacks based on gradient estimation spall2005introduction . Moon et al. moon2019parsimonious leveraged algorithms in combinatorial optimization. As opposed to white box gradient-based attacks, certain black-box adversarial attacks qiu2021black leverage Evolutionary Strategies to estimate the gradients used to perform a PGD-like attack. Alzantot et al. alzantot2019genattack recently reported about GenAttack, a gradient-free optimizer that uses Genetic Algorithms (GA) to produce subtle perturbations, which successfully fool state-of-the-art image recognition models with significantly fewer queries. Lu et al. FGBB propose to employ object detection techniques such as SIFT (Scale Invariant Feature Transform) to extract features from an image. The features are converted into a saliency distribution, where pixels that affect the composition of the image are given high probabilities. Then, crafting adversarial examples is framed as a strategic two-player game: one player manipulates features to minimize the distance to an adversarial example, while the other player adopts diverse strategies, i.e., cooperative, adversarial, or random. The resulting game states can be explored efficiently by Monte Carlo tree search to produce adversarial examples.
Many improvements have been proposed to upgrade white-box adversarial attacks for better transferability. SmoothFool (SF)s qiu2021black is an improved version of DeepFool that aims to produce smoother perturbations compared to the initial ones obtained through DeepFool. These smoother perturbations are intended to enhance the transferability of adversarial examples, making them more effective across different models or datasets. Dong et al. dong2018boosting developed an improved IFGSM with momentum to accumulate a velocity vector along the gradient direction. This accumulation of previous gradients stabilizes the updates and aids in navigating narrow valleys, small fluctuations, and suboptimal local extrema. Hence, the momentum-based IFGSM improves the transferability of the produced AXs across models. The momentum-based update is given by:
$v_{t+1}=\mu\cdot v_{t}+\frac{\nabla_{x}J(f(x_{t}),y)}{\|\nabla_{x}J(f(x_{t}),y)\|_{p}}$
The adversarial sample at the $(t+1)$-th step is computed as:
[…]
Here, $v_{t+1}$ is the updated velocity vector, $x_{t+1}$ is the adversarial sample at the $(t+1)$-th step, $x_{t}$ is the input at step $t$, $J$ represents the objective (cost) function, $f$ is the model, $y$ is the true label, $\alpha$ is the step size, $\epsilon$ is the maximum perturbation limit, $\mu$ is the momentum parameter, and $\|\cdot\|_{p}$ denotes the $L_{p}$ norm.
Inspired by meta-learning, Yuan et al. MGAA introduced a versatile method designed to enhance cross-model transferability within gradient-based attack frameworks. Their approach seamlessly integrates with any existing attack method and focuses on narrowing the disparity between white-box and black-box attack gradients. By randomly selecting models from a model zoo and iteratively conducting both white-box and black-box attacks on diverse tasks, they effectively enhance the transferability of adversarial examples across models in black-box scenarios.
@@
The first physical adversarial attack sharif2016accessorize was released in 2016 and it fools facial recognition systems by creating adversarial eyeglass frames designed with patterns that, when worn by individuals, cause mis-recognition of the individual. These patterns were strategically crafted to be imperceptible to humans while leading the facial recognition system to misidentify the wearer. This study marked an important milestone in the exploration of adversarial attacks in the physical domain, highlighting the susceptibility of facial recognition systems to imperceptible modifications applied to physical objects, such as glasses or eyeglass frames. Next, Engstrom et al. engstrom2017rotation showed the vulnerability of visual object recognition models against affine transformations (such as translations and rotations) of the images, which can be applied physically to the objects. More broader study gilmer2018motivating showed that visual object recognition models are prone to failing against simple guess-and-check of naturally-occurring situations related to the application domain, like taking pictures from another perspective angle.
Afterwards, more subtle physical attacks are proposed including patch-based and texture-based attacks. Patch-based attacks engender a universal adversarial image patch, which is stuck on the target object’s surface to mislead the DNNs. In performing patch-based physical attacks, the adversary prints the patch image and then sticks/hangs it on the surface of the target, covering its original appearance. For instance, a patch-based attack eykholt2018robust was proposed to deceive a traffic sign recognition model by generating strategically designed patches with specific patterns, colors, leading to erroneous predictions when placed on or near traffic signs. A more sophisticated patch-based attack wei2022adversarial inserts natural stickers that appear benign to humans but deceive facial recognition systems when placed on faces. Texture-based attacks manipulate the appearance of objects by generating adversarial textures applied to 3D models. The attacker first makes the adversarial texture physically, then wraps them over the target object’s surface, and the original texture is covered. For instance, the most representative form is the adversarial camouflage for vehicles wang2021dual , which subtly changes the visual appearance of vehicles, confusing detection models.
The environment-driven optical attacks exploit devices like projectors huang2022spaa , laser emitters duan2021adversarial , flashlights wang2023rfla , as well as natural occurrences like shadows zhong2022shadows and reflected light wang2023rfla . These methods are used to conduct physical adversarial attacks against visual models, taking advantage of environmental factors like lighting alterations, reflections, or other optical distortions that can negatively affect the model performance.
The emergence of generative models has expanded adversarial attacks to create meaningful perturbations aiming to generate semantic adversarial examples. For instance, SemanticAdv qiu2020semanticadv leverages attribute-conditioned image editing via a generator to alter specific semantic attributes like hair color, facial expressions, or a car’s position on the road. These modifications retain a realistic visual appearance and resemblance to the original image while changing specific semantic details.

### v3

@@ 1 Data Bias: Train-Serving Skew
The goal of classical ML and statistical learning theory is to achieve stable performance in an i.i.d. environment, the training data and the unseen data come from the same distribution bishop2006pattern . However, the i.i.d. assumption is not satisfied in most cases because the construction of training datasets with high probability to represent the true data distribution (the inputs in a production environment) is extremely difficult sugiyama2012machine . Real-world data is multi-faceted and virtually infinite, whereas training datasets are finite and constrained by the resources available during the dataset preparation. Data bias can occur in many ways, misrepresenting training datasets for real-world applications. This leads to a common failure known as the train-serving skew breck2017ml : models that perform well in development but poorly in deployment.
There are two main categories of data bias in ML applications: erroneous bias and discriminatory. Erroneous bias can be viewed as a systematic error caused by faulty assumptions. For instance, due to selection bias marlin2012collaborative or sampling bias mehrabi2021survey , the chosen training samples may not be able to represent the real data distribution.
Selection bias refers to a systematic error introduced in the data collection process where certain individuals are more likely to be selected than others, leading to a biased representation of the population.
Sampling bias occurs when the method used to retrieve data favors certain individuals or groups over others, resulting in a skewed or unrepresentative sample compared to the population of interest.
@@
Measurement bias happens when the device used to measure the signal has systematic value distortion that tends to skew the data in a particular direction that prevents the generalization of other data collected by other devices. Discriminatory bias reflects a model’s unfair behaviors toward a certain group or individual, such as producing discriminatory content or performing less well for a specific population or group shah2019predictive . The model is likely to inherit this discrimination bias from the training datasets, leading to undesirable performance on minority groups.
@@
Long-tailed distributions represent scenarios where certain classes or categories are significantly more prevalent than others, mirroring real-world data collection practices. These distributions are influenced by natural events, and not necessarily introduced by biased data collection. Nonetheless, ML algorithms fail to handle them properly as they are statistically optimized to perform well on common inputs (i.e. the head of the distribution) but struggle where examples are sparse (the tail). The tail often comprises the largest proportion of possible inputs, which makes their inclusion a laborious iterative procedure. Hence, ML models optimized on ID data following long-tail distributions are ineffective in dealing with edge cases because ERM relies on average cost minimization to learn statistically a model that is often biased towards the most common cases. Edge cases denote inputs lying at the extremes or boundaries of the expected range, often residing in the sparse tail of the data distribution. These cases are less frequently encountered in typical usage and common circumstances, but failing to consider them may result in unexpected behaviors and vulnerabilities in the ML model. In safety-critical domains such as self-driving cars and medical diagnosis, failure to handle edge cases properly may have catastrophic consequences, underscoring the importance of gathering and addressing them comprehensively brooks2017edge . Therefore, a robust ML workflow must address the challenges associated with long-tailed distribution modeling, especially considering edge cases that can have detrimental impacts in real-world use cases.
@@ 2 Model Complexity: A Double-Edge Sword
Conventionally, it is assumed that the use of models with increasing capacity will systematically result in overfitting the training data. Hence, the capacity of a model is usually controlled either by limiting the size of the model (number of parameters) or via various explicit or implicit regularizations, such as early stopping yao2007early , batch normalization ioffe2015batch , dropout srivastava2014dropout , and weight decay krogh1991simple . This aims to push learning to a subspace of a hypothesis with manageable complexity and reduce overfitting zhang2021understanding . Nonetheless, researchers have found that increasing model complexity not only allows for perfect interpolation but also results in low generalization error. Various studies have been conducted to analyze such overparameterized models, i.e., trainable parameters are much higher in number than the training data points. From the statistical viewpoint, the majority of overparameterized models exhibit a double-descent effect belkin2018understand ; mei2022generalization . In fact, the generalization error follows the traditional U-shaped curve until a specific point, after which the error decreases, and reaches a global minimum in the overparameterized regime. According to the double-descent phenomenon, the minimum generalization error tends to appear at infinite complexity, i.e., the more overparameterized the model, the smaller the error. From the optimization viewpoint, the good generalization behavior of highly overparameterized models is also commonly attributed to the inductive bias of gradient-based algorithms which helps with selecting models that generalize well despite the non-convexity, e.g., bartlett2021deep . Intuitively, the large number of hidden units here represent all possible features, and hence the optimization problem involves just picking the right features that will minimize the training loss. This suggests that as we over-parametrize the networks, the optimization algorithms need to do less work in tuning the weights of the hidden units to find the right solution.
@@
A disadvantage of overparameterized deep learning architectures is that they are highly susceptible to perturbations in adversarial or non-adversarial settings, compared to conventional, less sophisticated models. To illustrate the inherent brittleness of overparameterized neural networks, we refer to the notion of “neuron coverage” that is inspired by the code coverage in traditional software testing deepxplore . It involves generating synthetic test input data to trigger the neurons that have not been activated by the original test data. The success of this coverage criterion suggests that only a subset of the parameters is responsible for capturing the patterns needed for the task. The rest of the parameters might be unoptimized (almost stalled at initial random weights) or have received fewer updates over the training. The presence of these suboptimal subnetworks might not harm the model’s performance under i.i.d. conditions. However, it does affect the model’s robustness negatively. Any unusual changes in inputs that accidentally activate these neurons could lead to unpredictable model behaviors. In addition, any attacker can exploit the larger space of these suboptimal neurons by designing a malicious input that yields a particular model’s output.
@@ 3 Underspecified ML Pipeline: One Pipeline, Many Models
To solve an ML problem, we expect the model to encode some essential structure of the underlying distribution, which is inferred from the data using a designed ML pipeline, and is often what makes a model credible. Nevertheless, many explorations of the failures of ML pipelines that optimize for i.i.d. generalization, reveal a conflict between i.i.d. generalization and encoding credible inductive biases. It is called structural failure mode, as it is often diagnosed as a misalignment between the predictor learned by ERM and the causal structure of the desired predictor arjovsky2019invariant . In medical applications of ML, training inputs often include markers representing a doctor’s diagnostic judgments oakden2020hidden . An analysis of a CNN model for diagnosing skin lesions showed that it relied heavily on surgical ink markings around skin lesions that doctors had deemed cancerous winkler2019association . In these situations, a predictor with credible inductive biases cannot achieve optimal i.i.d. generalization in the training distribution, because there are so-called “spurious” features that are strongly associated with the label in the training data, but are not associated with the label in practically important settings. In fact, an iid-optimal predictor would incorporate the ink markings as they are highly correlated with positive cases, but these markings would not be expected to be present in deployment, where the predictor would itself be part of the workflow for making a diagnostic judgment. There is clearly an underspecification problem as there is not enough information (a lack of positivity) in the training distribution to distinguish between credible inductive biases and spurious relationships. Geirhos et al. shortcutLearning connect this underspecification issue to the notion of “shortcut learning”. They point out that there may be many predictors that generalize well in i.i.d. settings, but only a few of them align with the intended solution to the learning problem. Shortcut learning resembles surface learning of students in classrooms, relying on simple decision rules to pass an exam shortcutLearning . The problem with shortcuts is that they might go unnoticed during the i.i.d. testing and only occur in deployment scenarios. In the absence of large, diverse datasets, their risk is higher. For example, diagnostic data for rare or novel diseases is usually limited to small datasets, and unbiased validation data could be difficult to acquire.
@@
Modern ML pipelines are poorly set up for satisfying the system requirements underspecificationMLmodel . Their i.i.d. evaluation procedure often results in multiple models with equivalent (similar) predictive risk (performance) while they encode substantially different inductive biases. This implies that the ML pipeline could not distinguish between these iid-optimal models despite their potential differences in terms of robustness. The ML pipelines must be specified and designed in a way that promotes the selection of the model, encoding credible inductive biases, to bridge the gap between testing behavior and deployment behavior.
@@ 3 Robustness Assessment
1 Non-adversarial Shifts: From Synthetic to Real-World
Non-adversarial shifts, including natural data corruptions and perturbations, are rarely characterized. Only a few works in Computer Vision focus on such a type of model robustness assessment. They have introduced various benchmarking datasets carvalho2019machine to investigate the impact of naturally-shifted inputs on modern ML models. These benchmarks involve the introduction of corruptions and/or perturbations to standard (clean) datasets, allowing the assessment of different models’ resilience against corrupted images. For instance, ImageNet-C hendrycks2019benchmarking and ImageNet-P hendrycks2019benchmarking serve as synthetic benchmarks, each focusing on distinct aspects of robustness: corruption and perturbation. These benchmarks decouple robustness benchmarking by applying image transformations to the original images from the ImageNet dataset. Corruptions involve significant changes in image statistics, offering a testing ground for out-of-distribution scenarios. The benchmark hendrycks2019benchmarking includes 15 types of corruption transformations selected from noise, blur, weather, and digital categories, with five severity levels controlling the degree of distribution shift. The following are examples of algorithmically-generated corruptions:
    Gaussian noise simulates noise in low-lighting conditions.
    Shot (Poisson) noise represents fluctuations in pixel intensity.
    Impulse noise resembles sporadic bright or dark pixels.
    Defocus blur mimics the effect of an out-of-focus image.
    Frosted Glass Blur creates a soft, diffused effect.
    Motion blur represents blurring caused by rapid camera movement.
    Zoom blur occurs during rapid camera zooming.
    Snow visually obstructs details in an image.
    Frost creates a frosty or icy appearance.
    Fog creates a shrouding effect using the diamond-square algorithm.
    Brightness mimics changes in illumination intensity.
    Contrast levels vary depending on lighting conditions and object color.
    Pixelation occurs when upsampling a low-resolution image.
    JPEG compression introduces compression artifacts in images.
Indeed, each corruption varies in severity to reflect different levels of image degradation. For instance, speckle noise is an additive noise with severity controlled by the intensity of noise added to each pixel. Lower severity levels introduce minimal noise, while higher levels result in more noticeable distortions and reduced clarity. Similarly, Gaussian blur severity depends on the amount of blurring applied, determined by the size and weights of the filter. Lower severity levels apply mild blur, preserving most details, while higher levels lead to more pronounced blurring and loss of sharpness.
Perturbations, on the other hand, entail subtle transformations of original images, drawn from the same categories as corruptions but are designed to be more challenging to perceive visually. The perturbation benchmark hendrycks2019benchmarking aims to assess models’ performance in the face of subtle data distribution shifts. Creating a comprehensive real-world robustness benchmark that incorporates systematic distribution shifts poses challenges compared to synthetic benchmarks. The complexity arises from the multifaceted variations that can simultaneously occur in real-world data. For instance, images featuring the same object may be captured with different cameras, from diverse viewpoints, in various locations, and under different weather conditions recht2019imagenet ; hendrycks2021many . Temporal Perturbations shankar2021image are natural perturbations that are deduced from small changes occurring in assembled sets of contiguous video frames that appear perceptually similar to humans, but might produce inconsistent predictions for ML models. All these perturbations result in a condition where the distribution of the test set differs from the one of the training set.
Non-adversarial or natural perturbations serve as synthetic or newly created testing datasets for evaluating models in an out-of-distribution manner, simulating shifts from their original training distribution. These perturbations effectively highlight the models’ fragility to naturally occurring data distortions. However, the absence of performance degradation on these datasets does not necessarily imply robustness, as they may suffer from selection and sample bias. For instance, when creating datasets with data augmentation techniques like changing image brightness, the use of brightness-based corruptions might not be indicative of model robustness, as the training data distribution already encompasses various brightness levels. In this case, non-adversarial robustness represents how well the model maintains its predictive performance under foreseeable input shifts (i.e, variation of brightness level).
@@ 2 Adversarial Attacks: Categories and Aims
The predominant evaluation methods for ML model robustness are the adversarial attacks, where the community invents many ways to carefully craft perturbations that can deceive a given ML model. The purpose of such attacks is to produce an adversarial example (AX), i.e., an input $x^{\prime}$ close to a valid input $x$ according to some distance metric (i.e., similarity) or admissibility criteria (i.e., semantically-preserving modification ranges), whose model’s predictions, denoted as $f(x^{\prime})$ and $f(x)$, respectively, are different ($f(x^{\prime})\neq f(x)$). First, they can be categorized into white-box attacks and black-box attacks according to how much knowledge an attacker has about the subject model. White-box attacks are implemented with direct access to the model or its training data, whereas black-box attacks can only access the target model through queries: pairs of inputs and outputs. The growth of black-box attacks started with the discovery of the transferability of adversarial examples papernot2016transferability . Cross-model transferable, where attackers can construct adversarial examples in known deep learning models and subsequently attack a related unknown model. Cross training-set transferable refers to the attacks that exploit shared vulnerabilities across different datasets or domains. Second, adversarial attacks can be targeted or non-targeted. If we consider the example of image classification, targeted attacks aim to force the classifier to output a particular (chosen) class, whereas untargeted attacks attempt to make it return any class other than the original label. This categorization of adversarial attacks is based on the attacker’s goal and information access, which determines the relevance of an attack for an application. For instance, targeted white-box attacks are more suitable for security concerns, while untargeted black-box attacks can be appropriate for assessing robustness to noise.
@@
White-box adversarial attacks are a type of attack on ML models where the attacker has complete access to the architecture, parameters, and gradients of the target model. In this scenario, the attacker can exploit this knowledge on the model’s intrinsics to craft adversarial examples that are specifically designed to manipulate the model’s outcomes. Most white-box adversarial attacks employ Gradient-based approaches, where the attacker computes gradients of the model’s loss function with respect to the input data and uses this information to guide the perturbation process. The most notorious of these approaches is the Fast Gradient Sign Method (FGSM) FGSM . It operates as a one-step method by computing the gradient of the model’s cost function ($J$) with respect to the input data ($x$) and then perturbing the input along this gradient direction. Specifically, it alters the input ($x$) by adding noise ($\eta$) in the direction that maximizes the loss, using a magnitude defined by an epsilon value ($\epsilon$) to limit the perturbation within a certain range. The attack aims to maximize the loss by perturbing the input data without exceeding the epsilon-boundary, calculated using a distance metric (e.g., $L_{\infty}$ or $L_{2}$ norm). The formulation can be represented as:
@@
Despite its effectiveness in rapidly generating AXs, FGSM’s limitation lies in its single-step gradient-based approach to input perturbation. To address this, iterative variants of FGSM have been proposed. IFGSM IFGSM repeatedly applies FGSM steps with small size and clips pixel values, ensuring they are within an $\epsilon$-neighbourhood. PGD PGD is another iterative method, but randomly initializes examples within the $L_{\infty}$ norm sphere and performs random restarts, differing from IFGSM.
Aside from gradient-based methods, white-box adversarial attacks employ other techniques. Optimization-based approaches directly optimize perturbations to maximize the model’s prediction error while ensuring imperceptibility, such as Carlini & Wagner (C&W) carlini2017towards and DeepFool deepfool . Model-based approaches rely on a surrogate model to generate adversarial examples that transfer to the target model. For instance, AdvGAN AdvGAN employ GANs to generate adversarial examples by training the generator to craft perturbations that deceive the target model while optimizing the discriminator to differentiate between original and perturbed instances.
@@
Black-box adversarial attacks involve limited access to the target model’s architecture, parameters, and gradients. Despite this, attackers can interact with the model by querying input-output pairs to craft adversarial examples that alter its behaviors. Black-box adversarial attacks usually demand a high number of queries, but they are better representative of external attacks’ simulation.
To guide the search for adversarial perturbations through model queries, black-box adversaries may leverage local-search-based approaches narodytska2017simple , gradient estimation chen2017zoo ; bhagoji2018practical , combinatorial optimization moon2019parsimonious , and evolutionary metaheuristics qiu2021black like Genetic Algorithms (GA) alzantot2019genattack .
Many improvements have been proposed to upgrade white-box adversarial attacks for better transferability. SmoothFool (SF) smoothfool is an improved version of DeepFool that produces smoother perturbations to enhance the transferability of adversarial examples, making them more effective across different models or datasets. An IFGSM boosted with momentum dong2018boosting improves the transferability of the produced AXs across models as its stable updates aid in navigating narrow valleys, small fluctuations, and suboptimal local extrema.
Inspired by meta-learning, Meta Gradient Adversarial Attack (MGAA) MGAA is a versatile method designed to blur the boundary between white-box and black-box attack gradients. By randomly selecting models from a model zoo and iteratively conducting both white-box and black-box attacks on diverse tasks, MGAA effectively enhance the transferability of adversarial examples across models and seamlessly integrates with any existing attack method.
@@
The first physical adversarial attack sharif2016accessorize was released in 2016 and it fools facial recognition systems by creating adversarial eyeglass frames designed with patterns that, when worn by individuals, cause mis-recognition of the individual. It marked an important milestone in the exploration of adversarial attacks in the physical domain. Since then, many variants of physical adversarial attacks have been explored, each highlighting different concepts and vulnerabilities.
Physical attacks by object manipulation include affine transformations such as translations and rotations of the images that can be applied physically to the objects. They tend to expose brittleness of visual object recognition models. More broader, guess-and-check of naturally-occurring situations gilmer2018motivating , like changing the object position or taking pictures from another perspective angle, are able to fool visual object recognition models.
Patch- and texture-based attacks involve attaching adversarial patches or textures onto physical objects to deceive the targeted model. Patch-based attacks, as demonstrated by artificial sticks on traffic signs eykholt2018robust or natural stickers on faces wei2022adversarial , exploit strategically designed patches to be stuck on the target object’s surface. Texture-based attacks, as exemplified by adversarial camouflage for vehicleswang2021dual , manipulate the appearance of objects using adversarial textures applied to 3D models.
Environment-driven optical attacks manipulate maliciously the environmental factors like lighting alterations, reflections, or other optical distortions that can negatively affect the model performance. Examples include exploitation of devices like projectors huang2022spaa , laser emitters duan2021adversarial , flashlights wang2023rfla , as well as natural occurrences like shadows zhong2022shadows and reflected light wang2023rfla .
Model-based physical attacks leverage generative models to produce semantic adversarial examples. For instance, SemanticAdv qiu2020semanticadv leverages attribute-conditioned image editing via a generator to alter specific semantic attributes like hair color, facial expressions, or a car’s position on the road. These modifications retain a realistic visual appearance and resemblance to the original image while changing specific semantic details.